Differentiate f(x) = 5π‘₯Β³+7π‘₯ 2π‘₯Β²βˆ’4π‘₯+5 using quotient rule.Β 


1
Expert's answer
2021-08-23T07:48:26-0400

"f'(x)=\left(\frac{5x^3+7x}{2x^2-4x+5}\right)'=\\ {}=\frac{(5x^3+7x)'(2x^2-4x+5)-(5x^3+7x)(2x^2-4x+5)'}{(2x^2-4x+5)^2}=\\ {}=\frac{(15x^2+7)(2x^2-4x+5)-(5x^3+7x)(4x-4)}{(2x^2-4x+5)^2}=\\ {}=\frac{10x^4-40x^3+61x^2+35}{(2x^2-4x+5)^2}"